Calculate the change in entropy when argon at 25 degree Celcius and 1.00 bar in a container of volume 0.500 dm^3 is compressed to 0.500 dm^3 and cooled to 25 degree Celcius.
1
Expert's answer
2012-09-07T11:02:43-0400
Assuming ideal gas:
delta_S = n * Cv * ln(Tf / Ti) + n * R * ln(Vf / Vi)
argon molar heat capacity is 20.786 J·mol−1·K−
"when argon at 25 degree Celcius" and "oled to 25 degree Celcius" means that T is constant,the same with volume
delta_S = 1* 20.768* ln (1) + 1*8.31*ln (1) = 0 + 0 =0